A standard airbag has 130.0 g of sodium azide (NaN3).  How many grams of dangerous sodium metal would be produced? 


Show your stoichiometry work in the tables below, including all appropriate units, for full credit.  



1
Expert's answer
2021-04-22T03:03:57-0400

2 NaN3 → 2 Na + 3 N2

2 moles of NaN3 produce 2 moles of Na

Molar Mass of NaN3 = 65

= 130/65 = 2 moles of NaN3

= 2 moles of Na

Molar Mass of Na = 22.99

= 22.99 × 2

= 45.98g of Na